The decomposition of NI3 to form N2 and I2 releases −290.0 kJ of energy. The reaction can be represented as 2NI3(s)→N2(g)+3I2(g), ΔHrxn=−290.0 kJ Find the change in enthaply when 12.0 g of NI3 decomposes. Express your answer to three significant figures and include the appropriate units.
Answer #1

Answer:

Step 1: Explanation

When enthalpy is positive and delta H is greater than zero, this means that a system absorbed heat. This is called an endothermic reaction.

When enthalpy is negative and delta H is less than zero, this means that a system released heat. This is called an exothermic reaction

Step 2: Calculate the moles from mass

We know, moles = mass / molar mass

molar mass of NI3 = 394.719 g/mol

moles of NI3 = 12 g / 394.719 g/mol = 0.0304 mol

Step 2: Calculation of change in enthalpy due to heat released

Write the equation

2 NI3(s) --------->N2(g) + 3 I2(g); ΔH= -290 kJ

For 2 moles of NI3 enthalpy changed = -290 kJ

so, for 0.0304 mol of NI3 enthalpy change will be = ( -290 kJ / 2 mol ) × 0.0304 mol = -4.41 kJ